OpenStudy (anonymous):

The curve with the equation 2(x^2+y^2)^2=25(x^2-y^2) is called a lemniscate a) Find y' b) Find the equation of the tangent line to the lemniscate at the point (-3,1) c) Find the points on the lemniscate where the tangent line is horizontal.

OpenStudy (anonymous):

I got a) and b), but I don't know what to do for c). Please help! a) \[\frac{ dy }{ dx }=\frac{ x(25-4(x^2+y^2)) }{ y(25+4(x^2+y^2)) }\] b) 9x-13y+40=0

OpenStudy (zale101):

\(\large y'(x)=\LARGE\frac {-x(4x^2+4y^2-25)}{y(4x^2+4y^2+25)}\) to find all the points where the tangent is horizontal, you set the derivative numerator equal to zero but the denominator doesn't equal to zero to find all the points where the tangent is vertical, you set the derivative denominator equal to zero but the numerator doesn't equal to zero.

OpenStudy (zale101):

your question is asking for horizontal tangent. So solve that and substitute back to the original equation

OpenStudy (zale101):

solve for y, but since your original equation has \(2(x^2+y^2)^2=25(x^2-y^2)\) since the y's are y^2, make the equation equal to y^2 so you can easily substitute it back in

OpenStudy (anonymous):

I got this y^2=25/4+x^x? So I sub this equation to the original equation? @Zale101

OpenStudy (zale101):

Yes

OpenStudy (anonymous):

Opps i meant y^2=-25/4+x^2

OpenStudy (zale101):

and solve for x, after your solved for x, you'll get some number plug the number in x of this equation y^2=-25/4+x^2 and solve for y and from there, you'll have x and y (points)

OpenStudy (anonymous):

Is there an easier way to solve for this after subbing the equation? I got this for now: \[2(x^2+\frac{ 25 }{ 4 }+x^2)^2=25(x^2+\frac{ 25 }{ 4 }+x^2)\]

OpenStudy (zale101):

Well, i dont know of any easier method, but what you did looks alright to me

OpenStudy (anonymous):

I meant 2(x^2 - 25/4+x^2)^2...

OpenStudy (anonymous):

I got x=5/4?

OpenStudy (anonymous):

I think my answer is wrong :(

OpenStudy (zale101):

wolfram has something else :T

OpenStudy (anonymous):

My teacher told us that \[x=\pm \frac{ 5\sqrt{3} }{ 4 }\]

OpenStudy (anonymous):

can any of you guys help me with my math

OpenStudy (anonymous):

hummm...

OpenStudy (anonymous):

I can't get the answer

OpenStudy (zale101):

thisis what i keep getting \(−x(4x^2+4y^2−25)=0\) \(y^2=25-x^2 \) \(2(x^2+(\Large \frac{25}{4}-x^2))=25(x^2-(\Large\frac{25}{4}-x^2))\) \(x=\Large\frac{3\sqrt{\frac{3}{2}}}{2}\)

OpenStudy (anonymous):

Where did you get the negative sign from -x(4x^2+4y^2-25)=0?

OpenStudy (zale101):

\(2\frac{d}{dx}((x^2+y^2)^2=25\frac{d}{dx}(x^2-y^2)\) \((4x^2+4y^2)*(2x+2y*y')=25(2x-2y*y')\) simplify \(8x^3+8x^2y*y'+8xy^2+8y^3*y'=50x-50y*y'\) \(8x^2y*y'+8y^3*y'+50y*y=50x-8xy^2-8x^3\) \(8x^2y*y'+8y^3*y'+50y*y=-8x^3-8y^2x+50x\) \(y'=\Large\frac{-8x^3-8y^2x+50x}{8y^3+8x^2y+50y}\)

OpenStudy (zale101):

factor out a common factor, which is 2 \(y'=\Large\frac{-2x(4y^2+4x^2-25)}{2y(4y^2+4x^2+25)}=\Large\frac{-x(4y^2+4x^2-25)}{y(4y^2+4x^2+25)}\)

OpenStudy (anonymous):

Hmmm... The back of the book, there wasn't negative sign...

OpenStudy (zale101):

(shrugs) :/

OpenStudy (zale101):

i'm sure the implicit differentiation i did is correct. There must be something wrong....

OpenStudy (anonymous):

okay, thanks for helping me out! :) @Zale101